Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
SPB
SQS
XO
ToBeMigrated2XWiki
ToBeMigrated2GitlabPages
ToBeMigratedByGroup
Migrated2Zulip
Archived2PDF
not2MigrateOrArchive
SCS
SXP Migrated2Zulip
Photon Commissioning Team
XRO
DOC
HED Migrated2Zulip
WP74
MID
PSPO_XO
XO_Operation
SQS Migrated2Zulip
FXE Migrated2Zulip
DET
Sample Environment
SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness
Technical integration and coordination
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther
Not logged in
List
|
Find
|
Login
|
Help
Message ID:
32
Entry time:
12 Apr 2019, 10:41
Author:
Adriano Violante
Location:
General
Severity:
Information
Subject:
Status report after dispatch meeting 09.30
SPB/SFX
Hercules students period. During the night a pump failed and recovered after a power cycle.
5Hz mode tests were performed and successful
CSS being commissioned. X mov elmo stage still has to be configured. Will be done probably today.
After some updates, AGIPD run control topology light does not go to green. Start stop buttons do not work properly (cannot be seen when the detector stops aquiring data). CAS to have a look.
FXE
After some modifications in the chain fast valve-mps-plc a connector was found to be not properly reconnected and triggering the interlock. FXE should involve AE when such activities are performed in order not to cause overhead to the support groups.
SQS
SCS
Loop 4 updated. Loop 3 motors being commissioned. Loop 5 quadrant motion being finalized. Loop 12 (VAC) being prepared.
It is not clear when BKF 536 and 531 have to be installed by AE. SCS should comment.
MID
Loop 12 being prepared and should work by the end of next week.
HED
Loop 7 commissiong will start today. KS will send a list of crates to be put in the loops to AE (Prio-non prio PLCs)
OPT Lasers
SA2 PPL project, HW prepared.
XRO
XPD
VAC
DET
AE
ITDM
EET
CAS
SRP
TS
AoB
ELOG V3.1.4-7c3fd00